OpenStudy (anonymous):

Use the future value formula to compute the total amount of the investment. $4000 for 1 year at 5% compounded quarterly The amount in the account after 1 year is $_____

OpenStudy (amistre64):

sooo, what is the future value formula that youneed to use?

OpenStudy (anonymous):

I don't know

OpenStudy (amistre64):

well, they come in different flavors, so knowing which one its asking you to use would be helpful

OpenStudy (anonymous):

FV = PV(1+i)^n <- normal equation FV = PV(1+ (i/4))^n(4) <- quarterly so its FV = 4000( 1 + 0.05/4)^1(4) FV = 4000(1+ 0.0125)^4 FV = 4000(1.0125)^4 FV = 4000(1.05095) = $4203.78

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np glad to help